An issocles triagle had a perimeter of 370mm.the base of the triangle is 40mm longer than the two equal side.What is the length of each side.
Answer by checkley71(8403)

X+X+X+40=370
3X=370-40
3X=330
X=330/3
X=110 ANSWER FOR THE EQUAL SIDES.
PROOF
110+110+110+40=370
370=370